For each fixed λ > 0, let X have a Poisson distribution with parameter λ. Suppose λ


itself is a random variable with the gamma distribution


f(λ) =





1


Γ(n)


λ


n−1


e


−λ


, λ ≥ 0


0, λ < 0


where n is a fixed positive constant. Show that


P(X = k) = Γ(k + n)


Γ(n)Γ(k + 1) 


1


2


k+n


, k = 0, 1, 2

1
Expert's answer
2022-03-25T06:08:04-0400

"P\left( X=k \right) =\int_0^{+\infty}{P\left( X=k|\lambda =t \right) f_{\lambda}\left( t \right) dt}=\\=\int_0^{+\infty}{\frac{t^ke^{-t}}{k!}\frac{t^{n-1}e^{-t}}{\varGamma \left( n \right)}dt}=\\=\frac{1}{k!\varGamma \left( n \right)}\int_0^{+\infty}{t^{k+n-1}e^{-2t}dt}=\left[ 2t=x \right] =\\=\frac{1}{k!\varGamma \left( n \right)}\int_0^{+\infty}{\frac{1}{2^{k+n}}x^{k+n-1}e^{-x}dx}=\frac{\varGamma \left( k+n \right)}{\varGamma \left( k+1 \right) \varGamma \left( n \right) 2^{k+n}}"